OpenStudy (anonymous):

A hamster ball has a radius of 21 centimeters. What is its volume written in terms of π? Answer 12,348π cm3 9,261π cm3 6,946π cm3 588π cm3

OpenStudy (anonymous):

\[Volume =(4/3) \pi r^3\] so the volume = 12348 pi

jhonyy9 (jhonyy9):

(4/3)pir^3 = 12,348 cm^3 so the first

OpenStudy (anonymous):

\[Volume = {4 \over 3}\pi r^3 = { 4 \over 3 } \pi { (21)^3} = 12348 \pi cm^3\]